A round pizza, with diameter 48 cm, is into 15 equal sectors. A square pizza, with side length 44 cm, is cut into 25 equal squares. Which pizza slice, sector or square, has the greatest area? How much greater is it, to the nearest tenth of a square centimeter?
Answer by venugopalramana(3286)

A round pizza, with diameter 48 cm, is into 15 equal sectors.
TOTAL AREA PI*D^2/4=PI*48*48/4=576PI
EACH SECTOR AREA=576PI/15=192PI/5=120.576
A square pizza, with side length 44 cm, is cut into 25 equal squares.
TOTAL AREA OF SQUARE =44*44
EACH PIECE AREA=44*44/25=77.44
Which pizza slice, sector or square, has the greatest area? How much greater is it, to the nearest tenth of a square centimeter?
SO CIRCULAR PIZZA SLICE HAS GREATER AREA BY = 120.58-77.44=43.14 SQ.CM.